OpenStudy (anonymous):

A jeweler wants to make 1600 grams of 25% silver compound by mixing 20% and 40% silver compounds together. How many grams of each kind will she need?

OpenStudy (anonymous):

You need 1600 grams so x + y = 1600 .20x + .40y = .25(1600) x + y = 1600 20x + 40y = 25(1600) multiplied by 100 x + y = 1600 20x + 40y = 40000 -20x - 20y = -32000 multiplied by -20 20x + 40y = 40000 20y = 8000 y = 400 x + y = 1600 x + 400 = 1600 x = 1200 So 1200 grams of the 20% 400 grams of the 40%

OpenStudy (anonymous):

thanks